OpenStudy (anonymous):

compare the properties of the two functions problem.determine if the functions have a min or a max values and if their zeros are real or complex

OpenStudy (solomonzelman):

Hello:) and what are the functions?

OpenStudy (anonymous):

f(x)=25-2x-x^2 and g(x)=x^2-2x-25

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle f(x)=-x^2-2x+25 }\) \(\color{#000000}{ \displaystyle g(x)=x^2-2x-25 }\) I will re-write them that way.

OpenStudy (solomonzelman):

Well, g(x) has an \(x^2\) term in there. That means that the bigger the x gets (in this case, bigger=magnitude) the larger the output would get. Is there a limit to how large g(x) is going to get? (If there is, then, what)

OpenStudy (anonymous):

no sir dosent say a limit

OpenStudy (solomonzelman):

The bigger the x gets, the higher up will g(x) go. It will boundlessly go up. That means g(x) has no maximum.

OpenStudy (solomonzelman):

Although, g(x) would have a minimum. At the vertex.

OpenStudy (solomonzelman):

The rule is that if the parabola opens up (has a positive coefficient of x^2 term), then the minimum is the vertex, and no maximum. The rule is that if the parabola opens down (has a negative coefficient of x^2 term), then the maximum is the vertex, and no minimum.

OpenStudy (solomonzelman):

So, a function in a form: \(\color{#000000}{ \displaystyle h(x)= ax^2-bx+c }\) Will have, Minimum at the vertex. No Maximum. And a function in a form: \(\color{#000000}{ \displaystyle h(x)= -ax^2-bx+c }\) Will have, Maximum at the vertex. No Minimum.

OpenStudy (solomonzelman):

And to find the zeros, you will set each function =0, and solve for x.